Rio's company is manufacturing steel wire with an average tensile strength of 50 kilos. The Laboratory test 16 pieces showed that the mean was 45kls with standard deviation of 5 kilos , is it correct to the company to claim that the tensile strength of steel wire is 50 kilos? Test the hypothesis at 0.01 level of significance.

1
Expert's answer
2022-06-16T09:45:52-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=50"

"H_1:\mu\not=50"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=15" and the critical value for a two-tailed test is "t_c =2.946712."

The rejection region for this two-tailed test is "R = \{t:|t|>2.946712\}."

The t-statistic is computed as follows:



"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{45-50}{5/\sqrt{16}}=4"


Since it is observed that "|t|=4>2.946712=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=15" degrees of freedom, "t=4" is "p= 0.001159," and since "p= 0.001159<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 50, at the "\alpha = 0.01" significance level.